# What LLMs Miss in Recommendations: Bridging the Gap with Retrieval-Augmented Collaborative Signals

### v2

@@ 5. Conclusion
In the LLMs-as-recommender-systems (LLMs-as-RCs) paradigm, a key challenge is enabling LLMs to effectively incorporate collaborative information. In this paper, we began by analyzing the performance of LLMs in capturing collaborative signals for movie recommendation. We showed that a naive approach—embedding all user information directly into the prompt—makes it difficult for LLMs to interpret these signals, often performing worse than simple baselines such as matrix factorization.
To address this, we proposed a retrieval-augmented generation (RAG) based approach with improved prompting strategies. Our results demonstrate that presenting collaborative signals in a compact format and prompting the LLM to reason over them improves recommendation performance compared to traditional baselines.
Moreover, our method is both token-efficient and effective, achieving better results on standard evaluation metrics while minimizing prompt length.

### v3

@@ 5. Conclusion
In the LLMs-as-recommender-systems (LLMs-as-RCs) paradigm, a key challenge is enabling LLMs to effectively incorporate collaborative information. In this paper, we began by analyzing the performance of LLMs in capturing collaborative signals for movie recommendation. We showed that a naive approach—embedding all user information directly into the prompt—makes it difficult for LLMs to interpret these signals, often performing worse than simple baselines such as matrix factorization. To address this, we proposed a RAG based approach with improved prompting strategies. Our results demonstrate that presenting collaborative signals in a compact format and prompting the LLM to reason over them improves recommendation performance compared to traditional baselines. Moreover, our method is both token-efficient and effective, achieving better results on standard evaluation metrics while minimizing prompt length.